OpenStudy (anonymous):

solve equation, check solution r+2/5 = r-3/3

jigglypuff314 (jigglypuff314):

when you attempt to solve for r for r+2/5 = r-3/3 you would first try to get all r values to one side r+2/5 = r-3/3 -r -r but then you get 2/5 = -3/3 so no solution

OpenStudy (anonymous):

r + 2/5 = r -3/3 (subtract r from both sides) r - r + 2/5 = r - r -3/3 2/5 = 3/3 You will find this equation has no solution.